# GTA: A Benchmark for General Tool Agents

### v1

@@ C.1 Query Exemplars
We design several initial queries as query exemplars, as shown from Figure 22 to 36. The annotators brainstorm and design new questions that have the same tool chain as the exemplar but with different scenarios. We provide an expansion example for most exemplars for annotators to refer to.
@@ C.4 Illustration of Executable Tool Chains
An illustration on each part of the tool chain is shown in Figure 41. It is in the JSON format. It contains the involved tool list, file list, and dialog list. There are three roles in the dialog list: user, assistant, and tool. In the user’s dialog, the query content is recorded. In the assistant’s dialog, the correct tool call including the tool name and arguments is recorded. In the tool’s dialog, the tool’s return value is recorded. You can refer to Figure 7 to 12, Figure 13 to 16, and Figure 17 to 21 for JSON-format tool chain examples.
@@ D.2 ReAct-Style Prompts
The ReAct-style prompt template using for the agent system is shown in Figure 42. A prompt example is shown in Figure 43.
@@
[…]
Figure 43: A ReAct-style prompt example for the agent system.
D.3 Prediction Examples of Different Models
In order to visualize the tool-calling capabilities of each model, we provide some prediction examples of several models. The query of these predictions is shown in Figure 44. The prediction of different models are shown in Figure 45 to 59. We notice that models which have higher answer accuracy on GTA tend to give a longer tool chain, such as GPT-4, GPT-4o, Claude-3. LLaMA-3 tries to call tools but failed to follow a correct format. Some models tend to give a answer directly without calling tools, such as Qwen, Yi, Mistral.
Query: The men in the picture want to buy one NVIDIA GeForce RTX 4070 SUPER each. According to NVIDIA’s official website in January, how many dollars will they need to spend in total? Involved Tools: CountGivenObject, GoogleSearch, Calculator Files:
Steps: 1. Count the number of people in a photo. 2. Check the price of a RTX 4070 on NVIDIA’s website. 3. Calculate the total number of TX 4070 that need to be purchased. 4. Calculate the total price. Answer: 1797
Figure 44: A query example for different models.
{ "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "CountGivenObject", "arguments": { "image": "image/image_14.jpg", "text": "men" } } } ] }, { "role": "tool", "content": [ { "type": "text", "content": "3" } ] }, { "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "GoogleSearch", "arguments": { "query": "NVIDIA GeForce RTX 4070 SUPER price January site:nvidia.com", "k": 1 } } } ] },
Figure 45: (1 / 2) The prediction of GPT-4.
[…]
Figure 46: (2 / 2) The prediction of GPT-4.
{ "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "CountGivenObject", "arguments": { "image": "/home/PJLAB/mazerun/mmlab/opencompass/data/agentlego_bench_229/image/image_14.jpg", "text": "men" } } } ] }, { "role": "tool", "content": [ { "type": "text", "content": "3" } ] }, { "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "GoogleSearch", "arguments": { "query": "NVIDIA GeForce RTX 4070 SUPER price January 2023", "k": 1 } } } ] },
Figure 47: (1 / 2) The prediction of GPT-4o.
[…]
Figure 48: (2 / 2) The prediction of GPT-4o.
{ "role": "assistant", "tool_calls": [ { "type": "function", "function": { "name": "CountGivenObject", "arguments": { "image": "/home/PJLAB/mazerun/mmlab/opencompass/data/agentlego_bench_229/image/image_14.jpg", "text": "NVIDIA GeForce RTX 4070 SUPER" } } } ] }, { "role": "tool", "content": [ { "

### v2

@@ C.1 Query Exemplars
We design several initial queries as query exemplars, as shown from Figure 10 to 24. The annotators brainstorm and design new questions that have the same tool chain as the exemplar but with different scenarios. We provide an expansion example for most exemplars for annotators to refer to.
@@ C.4 Illustration of Executable Tool Chains
An illustration on each part of the tool chain is shown in Figure 29. It is in the JSON format. It contains the involved tool list, file list, and dialog list. There are three roles in the dialog list: user, assistant, and tool. In the user’s dialog, the query content is recorded. In the assistant’s dialog, the correct tool call including the tool name and arguments is recorded. In the tool’s dialog, the tool’s return value is recorded. You can refer to Figure 7 to LABEL:fig:obj_query_end, Figure 8 to LABEL:fig:sbj_query_end, and Figure 9 to LABEL:fig:img_query_end for JSON-format tool chain examples.
@@ D.2 ReAct-Style Prompts
The ReAct-style prompt template using for the agent system is shown in Figure 30.
@@
D.3 Final Answer Evaluation of Subjective and Image Generation Queries
For a subjective query, we use All-MPNet-Base-V2[25] to encode both the prediction and the ground truth. Then we calculate the cosine similarity between the two embeddings. To scale the score from 0 to 1, we consider only positive values:
$s=max\left(\frac{E_{pred}\cdot E_{gt}}{||E_{pred}||\cdot||E_{gt}||},0\right)$
For a query with image answers, the AnsAcc score can be formulated as:
$s=\prod_{i=1}^{n}I(t_{i}\in T_{pred})\cdot\text{SimScore}(arg_{i},arg_{pred})\in[0,1],$
where $t_{i}$, $arg_{i}$ is the $i$-th image generation-related tool (AddText, DrawBox, TextToImage, ImageStylization) in the ground truth tool chain. $I$ denotes the indicator function. $T_{preds}$ is the set of image generation-related tools that appears in the model prediction. $arg_{pred}$ is the corresponding argument for $t_{i}$ as predicted by the model.
D.4 Detailed Error Cases of Each Type
We conduct a detailed analysis of predictions generated by GPT-4-1106-Preview and Llama-3-8B-Instruct. We systematically categorize seven primary error types. Detailed error cases of each type are shown in Figure 31.
Figure 31: Detailed error cases of each type in the predictions generated by GPT-4-1106-Preview and Llama-3-8B-Instruct.
D.5 Comparison of Llama-2-Chat-7B and Agent-Flan-7B
We compare Llama-2-Chat-7B with Agent-Flan-7B on GTA benchmark to see if instruction tuning on ReAct and JSON format data can enhance the model’s performance. The comparison of the two models’ responses to a same user query is shown in Figure 32.
Figure 32: The comparison of Llama-2-Chat-7B and Agent-Flan-7B responses to a same user query.